# DiSCo Meets LLMs: A Unified Approach for Sparse Retrieval and Contextual Distillation in Conversational Search

### v1

@@ 5.3. Effectiveness-Efficiency Trade-off
In this subsection, we aim to answer RQ4 by analyzing the effecti-veness-efficiency trade-off of the proposed DiSCo models. First, we show the possibility of controlling the sparsity of the model, and thus efficiency. Then, we dive into sparsity at different depths of conversations.
@@
Controlling Sparsity. We focus here on the sparsity of the sparse representations, as a measure of efficiency within the inverted index. Figure 5 plots the performance and FLOPs of several DiSCo models when trained with different levels of sparsity. This is possible thanks to the regularization loss from the SPLADE architecture, controlled by the hyper-parameters (Formal et al., 2021b). From the Figure, we see for example that the same DiSCo T5 model can be trained with different degrees of regularization, leading to different FLOPs and MRR. Furthermore, Figure 5 shows the FLOPs of the associated teacher models: SPLADE MistralQR and T5QR. We notice that our student DiSCo models are more sparse compared to their teachers, as only having a constraint on the similarities during training allows more freedom on the representations and their sparsity. This answers part of RQ4, on the effect of the distillation on our model, as our DiSCo can be trained with different degrees of sparsity.
@@
Sparsity at different Conversation Depths. To further address RQ4, we plot the sparsity and effectiveness of several baselines in Figure 6. On the left part of the graph, we see the performance according to the depth of the conversation, and on the right the number of activated tokens in the representations. In particular, we see the difficulty for long conversations across models. Comparing our DiSCo with the convSPLADE baseline, we see that the main improvement comes from longer conversations. We also see that DiSCo outperforms its teacher model, SPLADE MistralQR, on the first few turns of the conversation. This is also a difficult task here as the TopiOCQA dataset has topic switches every few turns, making the context modeling task more complex.
@@
Looking at the sparsity, we can see that while conversations increase in depth, with lengths that go up to 350 tokens, representations of all models stay sparse, with only up to 55 non-zero dimensions. This shows that the models are able to learn the noise reduction task associated with the context modeling task. Now comparing the models, as SPLADE MistralQR relies on Mistral rewrite, we see that the representations are really consistent even for long conversations. On the other hand, DiSCo and convSPLADE sparsity decrease as the length of the conversation increases. DiSCo Mistral still shows better sparsity compared to ConvSPLADE.
It is also interesting to see that SPLADE MistralQR is both better and more sparse for higher depths, showing that it would be possible to have both a more efficient and effective model. This also answers RQ4, on the efficiency trade-off for longer conversations.
@@ 6. Conclusion & Future Work
In this work, we proposed to train sparse CS models with a relaxed distillation strategy unifying the two context modeling and retrieval tasks. Our experiments on LSR and DiSCo models demonstrate that this training objective brings important performance gains on both in-domain and out-of-domain retrieval. By distilling LLM rewrites, our method effectively learns from a single or several LLM teachers, and outperforms the teachers, reaching SoTA on several CS datasets. Since improving model efficiency is a key goal of distillation, we examine the sparsity and effectiveness of our approach. We see in particular that we can control the sparsity of the distilled models and that sparse representations can encode long conversations, without losing effectiveness. Furthermore, as our distillation process only requires scores, any model producing a similarity score could be distilled, including totally different architectures such as BM25 (Robertson, 2009) and cross-encoder models (Hofstätter et al., 2020) but we leave this for future work. Also, more studies could be made on improving the teachers model, as we limited ourselves to the 7B and 8B versions of Mistral and Llama. This could then increase the performance of the student. Finally, the distillation process could be applied for dense retrieval (Nogueira and Cho, 2019; Reimers and Gurevych, 2019), we leave this as future work.
@@ Acknowledgements.
This research was financed by the Swiss National Science Foundation, for the Personality And Conversational INformatiOn Access (PACINO) project. This funding is a collaboration between the University of Amsterdam and the Università della Svizzera italiana.References
@@
- Abbasiantaeb and Aliannejadi (2024) Zahra Abbasiantaeb and Mohammad Aliannejadi. 2024. Generate then Retrieve: Conversational Response Retrieval Using LLMs as Answer and Query Generators. arXiv preprint:2403.19302 (2024).
- Adlakha et al. (2022) Vaibhav Adlakha, Shehzaad Dhuliawala, Kaheer Suleman, Harm de Vries, and Siva Reddy. 2022. TopiOCQA: Open-domain Conversational Question Answering with Topic Switching. Transactions of the Association for Computational Linguistics 10 (2022), 468–483. https://doi.org/10.1162/tacl_a_00471
- Aliannejadi et al. (2024) Mohammad Aliannejadi, Zahra Abbasiantaeb, Shubham Chatterjee, Jeffrey Dalton, and Leif Azzopardi. 2024. TREC iKAT 2023: A Test Collection for Evaluating Conversational and Interactive Knowledge Assistants. In Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (Washington DC, USA) (SIGIR ’24). Association for Computing Machinery, New York, NY, USA, 819–829. https://doi.org/10.1145/3626772.3657860
- Anantha et al. (2021) Raviteja Anantha, Svitlana Vakulenko, Zhucheng Tu, Shayne Longpre, Stephen Pulman, and Srinivas Chappidi. 2021. Open-Domain Question Answering Goes Conversational via Question Rewriting. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Kristina Toutanova, Anna Rumshisky, Luke Zettlemoyer, Dilek Hakkani-Tur, Iz Beltagy, Steven Bethard, Ryan Cotterell, Tanmoy Chakraborty, and Yichao Zhou (Eds.). Association for Computational Linguistics, Online, 520–534. https://doi.org/10.18653/v1/2021.naacl-main.44
- Bassani (2022) Elias Bassani. 2022. ranx: A Blazing-Fast Python Library for Ranking Evaluation and Comparison. In ECIR (2) (Lecture Notes in Computer Science, Vol. 13186). Springer, 259–264. https://doi.org/10.1007/978-3-030-99739-7_30
- Cambazoglu et al. (2021) B. Barla Cambazoglu, Leila Tavakoli, Falk Scholer, Mark Sanderson, and Bruce Croft. 2021. An Intent Taxonomy for Questions Asked in Web Search. In Proceedings of the 2021 Conference on Human Information Interaction and Retrieval (Canberra ACT, Australia) (CHIIR ’21). Association for Computing Machinery, New York, NY, USA, 85–94. https://doi.org/10.1145/3406522.3446027
@@
- Dalton et al. (2022) Jeffrey Dalton, Sophie Fischer, Paul Owoicho, Filip Radlinski, Federico Rossetto, Johanne R. Trippas, and Hamed Zamani. 2022. Conversational Information Seeking: Theory and Application. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval (Madrid, Spain) (SIGIR ’22). Association for Computing Machinery, New York, NY, USA, 3455–3458. https://doi.org/10.1145/3477495.3532678
@@
- Dalton et al. (2020b) Jeffrey Dalton, Chenyan Xiong, Vaibhav Kumar, and Jamie Callan. 2020b. CAsT-19: A Dataset for Conversational Information Seeking. In Proceedings of the 43rd International ACM SIGIR Conference on Research and Development in Information Retrieval (Virtual Event, China) (SIGIR ’20). Association for Computing Machinery, New York, NY, USA, 1985–1988. https://doi.org/10.1145/3397271.3401206
@@
- Dubey et al. (2024) Abhimanyu Dubey, Abhinav Jauhri, Abhinav Pandey, Abhishek Kadian, Ahmad Al-Dahle, Aiesha Letman, Akhil Mathur, Alan Schelten, Amy Yang, Angela Fan, Anirudh Goyal, Anthony Hartshorn, Aobo Yang, Archi Mitra, Archie Sravankumar, Artem Korenev, Arthur Hinsvark, Arun Rao, Aston Zhang, Aurelien Rodriguez, Austen Gregerson, Ava Spataru, Baptiste Roziere, Bethany Biron, Binh Tang, Bobbie Chern, Charlotte Caucheteux, Chaya Nayak, Chloe Bi, Chris Marra, Chris McConnell, Christian Keller, Christophe Touret, Chunyang Wu, Corinne Wong, Cristian Canton Ferrer, Cyrus Nikolaidis, Damien Allonsius, Daniel Song, Danielle Pintz, Danny Livshits, David Esiobu, Dhruv Choudhary, Dhruv Mahajan, Diego Garcia-Olano, Diego Perino, Dieuwke Hupkes, Egor Lakomkin, Ehab AlBadawy, Elina Lobanova, Emily Dinan, Eric Michael Smith, Filip Radenovic, Frank Zhang, Gabriel Synnaeve, Gabrielle Lee, Georgia Lewis Anderson, Graeme Nail, Gregoire Mialon, Guan Pang, Guillem Cucurell, Hailey Nguyen, Hannah Korevaar, Hu Xu, Hugo Touvron, Iliyan Zarov, Imanol Arrieta Ibarra, Isabel Kloumann, Ishan Misra, Ivan Evtimov, Jade Copet, Jaewon Lee, Jan Geffert, Jana Vranes, Jason Park, Jay Mahadeokar, Jeet Shah, Jelmer van der Linde, Jennifer Billock, Jenny Hong, Jenya Lee, Jeremy Fu, Jianfeng Chi, Jianyu Huang, Jiawen Liu, Jie Wang, Jiecao Yu, Joanna Bitton, Joe Spisak, Jongsoo Park, Joseph Rocca, Joshua Johnstun, Joshua Saxe, Junteng Jia, Kalyan Vasuden Alwala, Kartikeya Upasani, Kate Plawiak, Ke Li, Kenneth Heafield, Kevin Stone, Khalid El-Arini, Krithika Iyer, Kshitiz Malik, Kuenley Chiu, Kunal Bhalla, Lauren Rantala-Yeary, Laurens van der Maaten, Lawrence Chen, Liang Tan, Liz Jenkins, Louis Martin, Lovish Madaan, Lubo Malo, Lukas Blecher, Lukas Landzaat, Luke de Oliveira, Madeline Muzzi, Mahesh Pasupuleti, Mannat Singh, Manohar Paluri, Marcin Kardas, Mathew Oldham, Mathieu Rita, Maya Pavlova, Melanie Kambadur, Mike Lewis, Min Si, Mitesh Kumar Singh, Mona Hassan, Naman Goyal, Narjes Torabi, Nikolay Bashlykov, Nikolay Bogoychev, Niladri Chatterji, Olivier Duchenne, Onur Çelebi, Patrick Alrassy, Pengchuan Zhang, Pengwei Li, Petar Vasic, Peter Weng, Prajjwal Bhargava, Pratik Dubal, Praveen Krishnan, Punit Singh Koura, Puxin Xu, Qing He, Qingxiao Dong, Ragavan Srinivasan, Raj Ganapathy, Ramon Calderer, Ricardo Silveira Cabral, Robert Stojnic, Roberta Raileanu, Rohit Girdhar, Rohit Patel, Romain Sauvestre, Ronnie Polidoro, Roshan Sumbaly, Ross Taylor, Ruan Silva, Rui Hou, Rui Wang, Saghar Hosseini, Sahana Chennabasappa, Sanjay Singh, Sean Bell, Seohyun Sonia Kim, Sergey Edunov, Shaoliang Nie, Sharan Narang, Sharath Raparthy, Sheng Shen, Shengye Wan, Shruti Bhosale, Shun Zhang, Simon Vandenhende, Soumya Batra, Spencer Whitman, Sten Sootla, Stephane Collot, Suchin Gururangan, Sydney Borodinsky, Tamar Herman, Tara Fowler, Tarek Sheasha, Thomas Georgiou, Thomas Scialom, Tobias Speckbacher, Todor Mihaylov, Tong Xiao, Ujjwal Karn, Vedanuj Goswami, Vibhor Gupta, Vignesh Ramanathan, Viktor Kerkez, Vincent Gonguet, Virginie Do, Vish Vogeti, Vladan Petrovic, Weiwei Chu, Wenhan Xiong, Wenyin Fu, Whitney Meers, Xavier Martinet, Xiaodong Wang, Xiaoqing Ellen Tan, Xinfeng Xie, Xuchao Jia, Xuewei Wang, Yaelle Goldschlag, Yashesh Gaur, Yasmine Babaei, Yi Wen, Yiwen Song, Yuchen Zhang, Yue Li, Yuning Mao, Zacharie Delpierre Coudert, Zheng Yan, Zhengxing Chen, Zoe Papakipos, Aaditya Singh, Aaron Grattafiori, Abha Jain, Adam Kelsey, Adam Shajnfeld, Adithya Gangidi, Adolfo Victoria, Ahuva Goldstand, Ajay Menon, Ajay Sharma, Alex Boesenberg, Alex Vaughan, Alexei Baevski, Allie Feinstein, Amanda Kallet, Amit Sangani, Anam Yunus, Andrei Lupu, Andres Alvarado, Andrew Caples, Andrew Gu, Andrew Ho, Andrew Poulton, Andrew Ryan, Ankit Ramchandani, Annie Franco, Aparajita Saraf, Arkabandhu Chowdhury, Ashley Gabriel, Ashwin Bharambe, Assaf Eisenman, Azadeh Yazdan, Beau James, Ben Maurer, Benjamin Leonhardi, Bernie Huang, Beth Loyd, Beto De Paola, Bhargavi Paranjape, Bing Liu, Bo Wu, Boyu Ni, Braden Hancock, Bram Wasti, Brandon Spence, Brani Stojkovic, Brian Gamido, Britt Montalvo, Carl Parker, Carly Burton, Catalina Mejia, Changhan Wang, Changkyu Kim, Chao Zhou, Chester Hu, Ching-Hsiang Chu, Chris Cai, Chris Tindal, Christoph Feichtenhofer, Damon Civin, Dana Beaty, Daniel Kreymer, Daniel Li, Danny Wyatt, David Adkins, David Xu, Davide Testuggine, Delia David, Devi Parikh, Diana Liskovich, Didem Foss, Dingkang Wang, Duc Le, Dustin Holland, Edward Dowling, Eissa Jamil, Elaine Montgomery, Eleonora Presani, Emily Hahn, Emily Wood, Erik Brinkman, Esteban Arcaute, Evan Dunbar, Evan Smothers, Fei Sun, Felix Kreuk, Feng Tian, Firat Ozgenel, Francesco Caggioni, Francisco Guzmán, Frank Kanayet, Frank Seide, Gabriela Medina Florez, Gabriella Schwarz, Gada Badeer, Georgia Swee, Gil Halpern, Govind Thattai, Grant Herman, Grigory Sizov, Guangyi, Zhang, Guna Lakshminarayanan, Hamid Shojanazeri, Han Zou, Hannah Wang, Hanwen Zha, Haroun Habeeb, Harrison Rudolph, Helen Suk, Henry Aspegren, Hunter Goldman, Ibrahim Damlaj, Igor Molybog, Igor Tufanov, Irina-Elena Veliche, Itai Gat, Jake Weissman, James Geboski, James Kohli, Japhet Asher, Jean-Baptiste Gaya, Jeff Marcus, Jeff Tang, Jennifer Chan, Jenny Zhen, Jeremy Reizenstein, Jeremy Teboul, Jessica Zhong, Jian Jin, Jingyi Yang, Joe Cummings, Jon Carvill, Jon Shepard, Jonathan McPhie, Jonathan Torres, Josh Ginsburg, Junjie Wang, Kai Wu, Kam Hou U, Karan Saxena, Karthik Prasad, Kartikay Khandelwal, Katayoun Zand, Kathy Matosich, Kaushik Veeraraghavan, Kelly Michelena, Keqian Li, Kun Huang, Kunal Chawla, Kushal Lakhotia, Kyle Huang, Lailin Chen, Lakshya Garg, Lavender A, Leandro Silva, Lee Bell, Lei Zhang, Liangpeng Guo, Licheng Yu, Liron Moshkovich, Luca Wehrstedt, Madian Khabsa, Manav Avalani, Manish Bhatt, Maria Tsimpoukelli, Martynas Mankus, Matan Hasson, Matthew Lennie, Matthias Reso, Maxim Groshev, Maxim Naumov, Maya Lathi, Meghan Keneally, Michael L. Seltzer, Michal Valko, Michelle Restrepo, Mihir Patel, Mik Vyatskov, Mikayel Samvelyan, Mike Clark, Mike Macey, Mike Wang, Miquel Jubert Hermoso, Mo Metanat, Mohammad Rastegari, Munish Bansal, Nandhini Santhanam, Natascha Parks, Natasha White, Navyata Bawa, Nayan Singhal, Nick Egebo, Nicolas Usunier, Nikolay Pavlovich Laptev, Ning Dong, Ning Zhang, Norman Cheng, Oleg Chernoguz, Olivia Hart, Omkar Salpekar, Ozlem Kalinli, Parkin Kent, Parth Parekh, Paul Saab, Pavan Balaji, Pedro Rittner, Philip Bontrager, Pierre Roux, Piotr Dollar, Polina Zvyagina, Prashant Ratanchandani, Pritish Yuvraj, Qian Liang, Rachad Alao, Rachel Rodriguez, Rafi Ayub, Raghotham Murthy, Raghu Nayani, Rahul Mitra, Raymond Li, Rebekkah Hogan, Robin Battey, Rocky Wang, Rohan Maheswari, Russ Howes, Ruty Rinott, Sai Jayesh Bondu, Samyak Datta, Sara Chugh, Sara Hunt, Sargun Dhillon, Sasha Sidorov, Satadru Pan, Saurabh Verma, Seiji Yamamoto, Sharadh Ramaswamy, Shaun Lindsay, Shaun Lindsay, Sheng Feng, Shenghao Lin, Shengxin Cindy Zha, Shiva Shankar, Shuqiang Zhang, Shuqiang Zhang, Sinong Wang, Sneha Agarwal, Soji Sajuyigbe, Soumith Chintala, Stephanie Max, Stephen Chen, Steve Kehoe, Steve Satterfield, Sudarshan Govindaprasad, Sumit Gupta, Sungmin Cho, Sunny Virk, Suraj Subramanian, Sy Choudhury, Sydney Goldman, Tal Remez, Tamar Glaser, Tamara Best, Thilo Kohler, Thomas Robinson, Tianhe Li, Tianjun Zhang, Tim Matthews, Timothy Chou, Tzook Shaked, Varun Vontimitta, Victoria Ajayi, Victoria Montanez, Vijai Mohan, Vinay Satish Kumar, Vishal Mangla, Vítor Albiero, Vlad Ionescu, Vlad Poenaru, Vlad Tiberiu Mihailescu, Vladimir Ivanov, Wei Li, Wenchen Wang, Wenwen Jiang, Wes Bouaziz, Will Constable, Xiaocheng Tang, Xiaofang Wang, Xiaojian Wu, Xiaolan Wang, Xide Xia, Xilun Wu, Xinbo Gao, Yanjun Chen, Ye Hu, Ye Jia, Ye Qi, Yenda Li, Yilin Zhang, Ying Zhang, Yossi Adi, Youngjin Nam, Yu, Wang, Yuchen Hao, Yundi Qian, Yuzi He, Zach Rait, Zachary DeVito, Zef Rosnbrick, Zhaoduo Wen, Zhenyu Yang, and Zhiwei Zhao. 2024. The Llama 3 Herd of Models. arXiv preprint:2407.21783 (2024).
@@
- Elgohary et al. (2019) Ahmed Elgohary, Denis Peskov, and Jordan Boyd-Graber. 2019. Can You Unpack That? Learning to Rewrite Questions-in-Context. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), Kentaro Inui, Jing Jiang, Vincent Ng, and Xiaojun Wan (Eds.). Association for Computational Linguistics, Hong Kong, China, 5918–5924. https://doi.org/10.18653/v1/D19-1605
@@
- Formal et al. (2022) Thibault Formal, Carlos Lassance, Benjamin Piwowarski, and Stéphane Clinchant. 2022. From Distillation to Hard Negative Sampling: Making Sparse Neural IR Models More Effective. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval (Madrid, Spain) (SIGIR ’22). Association for Computing Machinery, New York, NY, USA, 2353–2359. https://doi.org/10.1145/3477495.3531857
- Formal et al. (2021b) Thibault Formal, Benjamin Piwowarski, and Stéphane Clinchant. 2021b. SPLADE: Sparse Lexical and Expansion Model for First Stage Ranking. In Proceedings of the 44th International ACM SIGIR Conference on Research and Development in Information Retrieval (Virtual Event, Canada) (SIGIR ’21). Association for Computing Machinery, New York, NY, USA, 2288–2292. https://doi.org/10.1145/3404835.3463098
@@
- Hai Le et al. (2023) Nam Hai Le, Thomas Gerald, Thibault Formal, Jian-Yun Nie, Benjamin Piwowarski, and Laure Soulier. 2023. CoSPLADE: Contextualizing SPLADE for Conversational Information Retrieval. In Advances in Information Retrieval: 45th European Conference on Information Retrieval, ECIR 2023, Dublin, Ireland, April 2–6, 2023, Proceedings, Part I (Dublin, Ireland). Springer-Verlag, Berlin, Heidelberg, 537–552. https://doi.org/10.1007/978-3-031-28244-7_34
@@
- Jang et al. (2024) Yunah Jang, Kang-il Lee, Hyunkyung Bae, Hwanhee Lee, and Kyomin Jung. 2024. IterCQR: Iterative Conversational Query Reformulation with Retrieval Guidance. In Proceedings of the 2024 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), Kevin Duh, Helena Gomez, and Steven Bethard (Eds.). Association for Computational Linguistics, Mexico City, Mexico, 8121–8138. https://doi.org/10.18653/v1/2024.naacl-long.449
- Järvelin and Kekäläinen (2000) Kalervo Järvelin and Jaana Kekäläinen. 2000. IR Evaluation Methods for Retrieving Highly Relevant Documents. In Proceedings of the 23rd Annual International ACM SIGIR Conference on Research and Development in Information Retrieval (Athens, Greece) (SIGIR ’00). ACM, New York, NY, USA, 41–48. https://doi.org/10.1145/345508.345545
@@
- Lassance and Clinchant (2022a) Carlos Lassance and Stéphane Clinchant. 2022a. An Efficiency Study for SPLADE Models. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval (Madrid, Spain) (SIGIR ’22). Association for Computing Machinery, New York, NY, USA, 2220–2226. https://doi.org/10.1145/3477495.3531833

### v2

@@ 5.3. Effectiveness-Efficiency Trade-off
In this subsection, we aim to answer RQ4 by analyzing the effecti-veness-efficiency trade-off of the proposed DiSCo models. First, we measure the efficiency of our method compared to rewriting-based methods. Then, we show the possibility of controlling the sparsity of the model through the FLOP regularization. Finally, we dive into sparsity at different depths of conversations.
Efficiency. As to explicit the inference efficiency gain of our method compared to rewrite-based approaches, we plot in Table 5 the efficiency in milliseconds of the models on TREC CAsT 2020. We observe from the table that while SPLADE MistralQR achieves very high performance here out-of-domain, the efficiency is very low compared to DiSCo Mistral. This is because of the rewriting step, which involves 1 LLM call for MistralQR. Also note that the rewrite needs to be executed on GPU, while most of the retrieval time is a search through the inverted index on CPU. DiSCo Mistral and multi-teach thus appear to have a better trade-off when it comes to efficiency. Finally, DiSCo Fusion has the best performance overall, without an important computational overload compared to SPLADE MistralQR 555DiSCo Fusion being an ensembling of two approaches, we report the maximum execution time of the two (as they could be executed in parallel).. Note that retrieval is not optimized, and libraries such as PISA could reduce retrieval of SPLADE models below 100 ms on CPU, as shown in recent papers (Lassance and Clinchant, 2022a; Lassance et al., 2023, 2024a). This would further increase the efficiency gap between rewrite-based and distillation-based approaches. We also focus here on inference efficiency, as training efficiency does not directly affect the user, and would involve training LLMs. Those results provide a first answer to RQ4 on the efficiency of DiSCo compared to existing methods.
@@
Controlling Sparsity. We then focus here on the sparsity of the sparse representations, as a measure of efficiency within the inverted index. Figure 5 plots the performance and FLOPs of several DiSCo models when trained with different levels of sparsity. This is possible thanks to the regularization loss from the SPLADE architecture, controlled by the hyper-parameters (Formal et al., 2021b). From the Figure, we see for example that the same DiSCo T5 model can be trained with different degrees of regularization, leading to different FLOPs and MRR. Furthermore, Figure 5 shows the FLOPs of the associated teacher models: SPLADE MistralQR and T5QR. We notice that our student DiSCo models are more sparse compared to their teachers, as having only a constraint on similarities during training allows more freedom on the representations and their sparsity. This answers part of RQ4, showing that DiSCo can be trained with different sparsity levels.
@@
Sparsity at different Conversation Depths. To further address RQ4, we plot the sparsity and effectiveness of several baselines in Figure 6. On the left part of the graph, we see the performance according to the depth of the conversation – as the number of previous interactions between the user and the system – and on the right the number of activated tokens in the representations. In particular, we see the difficulty for long conversations across models. Comparing our DiSCo with the convSPLADE baseline, we see that the main improvement comes from longer conversations. We also see that DiSCo outperforms its teacher model, SPLADE MistralQR, on the first few turns of the conversation. This is also a difficult task here as the TopiOCQA dataset has topic switches every few turns, making the context modeling task more complex.
@@
Examining sparsity, we observe that the representations across all models remain sparse, with at most 60 non-zero dimensions even at deeper conversation levels (). At these depths, conversations become significantly longer, reaching up to 350 tokens. This suggests that the models effectively perform noise reduction as part of the context modeling task, activating only a subset of the tokens from the context. Now, comparing the models, as SPLADE MistralQR relies on Mistral rewrite, we see that the representation sparsity is consistent even for long conversations. For DiSCo and convSPLADE, sparsity decreases with the turns of the conversation.
@@ 6. Conclusion & Future Work
In this work, we propose DiSCo, a novel distillation strategy in CS to distill query rewriting from LLMs. DiSCo trains CS models with a relaxed distillation strategy that unifies context modeling and retrieval tasks. Our experiments on LSR and DiSCo models demonstrate that this training objective achieves important performance gains in both in-domain and out-of-domain retrieval. By distilling LLM rewrites, our method effectively learns from a single or several LLM teachers, outperforming the teachers, and reaching SoTA on several CS datasets. The proposed training strategy of DiSCo also shows robustness to the quality of the teacher model when trained on weaker teacher models. We further examine the inference efficiency and sparsity of our approach after distillation. Our findings emphasize the importance of aligning the rewriting and retrieval tasks in CS, with training objectives that unify both tasks. As a future work, we plan to study different teachers, as any model producing a similarity score could be distilled, e.g., cross-encoder (Hofstätter et al., 2020), paired with stronger LLMs for rewrite.
@@ Acknowledgements.
This research was partly supported by the Swiss National Science Foundation (SNSF), under the project PACINO (Personality And Conversational INformatiOn Access), grant number 215742. This funding is a collaboration between the University of Amsterdam and the Università della Svizzera italiana.References
@@
- Abbasiantaeb et al. (2024) Zahra Abbasiantaeb, Simon Lupart, and Mohammad Aliannejadi. 2024. Generating Multi-Aspect Queries for Conversational Search. arXiv preprint arXiv:2403.19302 (2024).
- Adlakha et al. (2022) Vaibhav Adlakha, Shehzaad Dhuliawala, Kaheer Suleman, Harm de Vries, and Siva Reddy. 2022. TopiOCQA: Open-domain Conversational Question Answering with Topic Switching. Transactions of the Association for Computational Linguistics 10 (2022), 468–483. doi:10.1162/tacl_a_00471
- Aliannejadi et al. (2024) Mohammad Aliannejadi, Zahra Abbasiantaeb, Shubham Chatterjee, Jeffrey Dalton, and Leif Azzopardi. 2024. TREC iKAT 2023: A Test Collection for Evaluating Conversational and Interactive Knowledge Assistants. In Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (Washington DC, USA) (SIGIR ’24). Association for Computing Machinery, New York, NY, USA, 819–829. doi:10.1145/3626772.3657860
- Anantha et al. (2021) Raviteja Anantha, Svitlana Vakulenko, Zhucheng Tu, Shayne Longpre, Stephen Pulman, and Srinivas Chappidi. 2021. Open-Domain Question Answering Goes Conversational via Question Rewriting. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Kristina Toutanova, Anna Rumshisky, Luke Zettlemoyer, Dilek Hakkani-Tur, Iz Beltagy, Steven Bethard, Ryan Cotterell, Tanmoy Chakraborty, and Yichao Zhou (Eds.). Association for Computational Linguistics, Online, 520–534. doi:10.18653/v1/2021.naacl-main.44
- Bassani (2022) Elias Bassani. 2022. ranx: A Blazing-Fast Python Library for Ranking Evaluation and Comparison. In ECIR (2) (Lecture Notes in Computer Science, Vol. 13186). Springer, 259–264. doi:10.1007/978-3-030-99739-7_30
- Cambazoglu et al. (2021) B. Barla Cambazoglu, Leila Tavakoli, Falk Scholer, Mark Sanderson, and Bruce Croft. 2021. An Intent Taxonomy for Questions Asked in Web Search. In Proceedings of the 2021 Conference on Human Information Interaction and Retrieval (Canberra ACT, Australia) (CHIIR ’21). Association for Computing Machinery, New York, NY, USA, 85–94. doi:10.1145/3406522.3446027
@@
- Dalton et al. (2022) Jeffrey Dalton, Sophie Fischer, Paul Owoicho, Filip Radlinski, Federico Rossetto, Johanne R. Trippas, and Hamed Zamani. 2022. Conversational Information Seeking: Theory and Application. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval (Madrid, Spain) (SIGIR ’22). Association for Computing Machinery, New York, NY, USA, 3455–3458. doi:10.1145/3477495.3532678
@@
- Dalton et al. (2020b) Jeffrey Dalton, Chenyan Xiong, Vaibhav Kumar, and Jamie Callan. 2020b. CAsT-19: A Dataset for Conversational Information Seeking. In Proceedings of the 43rd International ACM SIGIR Conference on Research and Development in Information Retrieval (Virtual Event, China) (SIGIR ’20). Association for Computing Machinery, New York, NY, USA, 1985–1988. doi:10.1145/3397271.3401206
@@
- Dubey et al. (2024) Abhimanyu Dubey, Abhinav Jauhri, Abhinav Pandey, Abhishek Kadian, Ahmad Al-Dahle, Aiesha Letman, Akhil Mathur, Alan Schelten, Amy Yang, Angela Fan, Anirudh Goyal, Anthony Hartshorn, and et al. 2024. The Llama 3 Herd of Models. CoRR abs/2407.21783 (2024). doi:10.48550/ARXIV.2407.21783 arXiv:2407.21783
@@
- Elgohary et al. (2019) Ahmed Elgohary, Denis Peskov, and Jordan Boyd-Graber. 2019. Can You Unpack That? Learning to Rewrite Questions-in-Context. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), Kentaro Inui, Jing Jiang, Vincent Ng, and Xiaojun Wan (Eds.). Association for Computational Linguistics, Hong Kong, China, 5918–5924. doi:10.18653/v1/D19-1605
@@
- Formal et al. (2022) Thibault Formal, Carlos Lassance, Benjamin Piwowarski, and Stéphane Clinchant. 2022. From Distillation to Hard Negative Sampling: Making Sparse Neural IR Models More Effective. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval (Madrid, Spain) (SIGIR ’22). Association for Computing Machinery, New York, NY, USA, 2353–2359. doi:10.1145/3477495.3531857
- Formal et al. (2021b) Thibault Formal, Benjamin Piwowarski, and Stéphane Clinchant. 2021b. SPLADE: Sparse Lexical and Expansion Model for First Stage Ranking. In Proceedings of the 44th International ACM SIGIR Conference on Research and Development in Information Retrieval (Virtual Event, Canada) (SIGIR ’21). Association for Computing Machinery, New York, NY, USA, 2288–2292. doi:10.1145/3404835.3463098
@@
- Hai Le et al. (2023) Nam Hai Le, Thomas Gerald, Thibault Formal, Jian-Yun Nie, Benjamin Piwowarski, and Laure Soulier. 2023. CoSPLADE: Contextualizing SPLADE for Conversational Information Retrieval. In Advances in Information Retrieval: 45th European Conference on Information Retrieval, ECIR 2023, Dublin, Ireland, April 2–6, 2023, Proceedings, Part I (Dublin, Ireland). Springer-Verlag, Berlin, Heidelberg, 537–552. doi:10.1007/978-3-031-28244-7_34
@@
- Jang et al. (2024) Yunah Jang, Kang-il Lee, Hyunkyung Bae, Hwanhee Lee, and Kyomin Jung. 2024. IterCQR: Iterative Conversational Query Reformulation with Retrieval Guidance. In Proceedings of the 2024 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), Kevin Duh, Helena Gomez, and Steven Bethard (Eds.). Association for Computational Linguistics, Mexico City, Mexico, 8121–8138. doi:10.18653/v1/2024.naacl-long.449
- Järvelin and Kekäläinen (2000) Kalervo Järvelin and Jaana Kekäläinen. 2000. IR Evaluation Methods for Retrieving Highly Relevant Documents. In Proceedings of the 23rd Annual International ACM SIGIR Conference on Research and Development in Information Retrieval (Athens, Greece) (SIGIR ’00). ACM, New York, NY, USA, 41–48. doi:10.1145/345508.345545
@@
- Lassance and Clinchant (2022a) Carlos Lassance and Stéphane Clinchant. 2022a. An Efficiency Study for SPLADE Models. In Proceedings of the 45th International ACM SIGIR Conference on Research and Development in Information Retrieval (Madrid, Spain) (SIGIR ’22). Association for Computing Machinery, New York, NY, USA, 2220–2226. doi:10.1145/3477495.3531833